1. A drinking glass is in the shape of a frustum of a cone of height 14 cm. The diameters of its two circular ends are 4 cm and 2 cm. Find the capacity of the glass.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Radius (r1) of the upper base = 4/2 = 2 cm Radius (r2) of lower the base = 2/2 = 1 cm Height = 14 cm Now, Capacity of glass = Volume of frustum of cone So, Capacity of glass = (⅓)×π×h(r12+r22+r1r2) = (⅓)×π×(14)(22+12+ (2)(1)) ∴ The capacity of the glass = 102×(⅔) cm3